Which fraction is the greatest: 15/30, 4/15, 3/48, or 12/36?

Answer by Edwin McCravy(20056)

15/30, 4/15, 3/48, or 12/36

If you have a calculator, you can easily tell
their relative size by dividing numerator by denominator:

15/30 = 0.5
 4/15 = 0.26666666...
 3/48 = 0.0625
12/36 = 0.33333333...

The one with the largest tenths digits is 15/30.  

If you cannot use a calculator, reduce them all to
lowest terms:

15/30, 4/15, 3/48, 12/36

  1/2, 4/15, 1/16,  1/3

Get their LCM (or LCD) of 240

120/240, 64/240, 15/240, 80/240 

And the one with the largest numerator is 120/240
is 15/30.

Edwin